# ADEdgeDrop: Adversarial Edge Dropping for Robust Graph Neural Networks

### v1

@@ V-B Experimental Results
TABLE IV: Ablation study of ADEdgeDrop with GCN as the backbone. ADEdgeDrop w/o AT: ADEdgeDrop without the Adversarial Training.
Datasets ADEdgeDrop w/o AT ADEdgeDrop
Citeseer 0.708 0.721
Pubmed 0.791 0.794
ACM 0.886 0.893
Chameleon 0.504 0.523
@@ VI Conclusion
In this paper, we proposed a novel edge-dropping strategy named ADEdgeDrop, which integrates the downstream GNNs with an adversarial edge predictor supervising the edge-dropping procedure in contrast to conventional random edge-dropping methods. We optimized the edge predictor by formulating a saddle point problem to seek the most vicious perturbations so that robust optimizations of both the edge predictor and the basic GNNs can be achieved. We guided training algorithms of perturbations and other weights by PGD and SGD to guarantee the full optimization of all trainable variables. Comprehensive experiments on six benchmark datasets demonstrated that the proposed approach outperformed the other state-of-the-art GAL baselines. Moreover, ADEdgeDrop succeeded in improving the robustness of GNNs under different types of edge attacks. In future work, we will further investigate robust GAL methods with the simultaneous addition and removal of edges.

### v2

@@ VI Conclusion
In this paper, we have proposed a novel edge-dropping strategy named ADEdgeDrop, which integrates the downstream GNNs with an adversarial edge predictor supervising the edge-dropping procedure in contrast to conventional random edge-dropping methods. We optimized the edge predictor by formulating a saddle point problem to seek the most vicious perturbations so that robust optimizations of both the edge predictor and the basic GNNs can be achieved. We guided training algorithms of perturbations and other weights by PGD and SGD to guarantee the full optimization of all trainable variables. Comprehensive experiments on eight benchmark datasets have demonstrated that the proposed approach outperforms the state-of-the-art GAL baselines. Moreover, ADEdgeDrop succeeded in improving the robustness of GNNs under different types of edge-removal or edge-addition attacks. In future work, we will further investigate robust GAL methods with the simultaneous addition and removal of edges.